OpenStudy (anonymous):

Eigenvector calculation, i have worked out everything for you! just need verification!

OpenStudy (anonymous):

M= 2 0 -2 0 4 0 -2 0 5 eigenvalues are 1, 4, 6, need eigenvector for eigenvalue = 4 please!!

OpenStudy (amistre64):

subtract 4 from the diags and rref

OpenStudy (anonymous):

i have done that but you get M- lambda = -2 0 -2 0 0 0 -2 0 -1 !!!

OpenStudy (amistre64):

keep reducing ...

OpenStudy (anonymous):

M3,3 = 1 not -1... sorry

OpenStudy (amistre64):

-2 0 -2 0 0 0 -2 0 1 1 0 1 0 0 0 -1 0 1/2 1 0 1 0 0 0 0 0 3/2 1 0 1 0 0 0 0 0 1 1 0 0 0 0 0 0 0 1 right?

OpenStudy (anonymous):

which implies eigen vector of (0,0,0)!!! can this be possible? or am i wrong?

OpenStudy (amistre64):

x1 0 x2 = x2 1 x3 0

OpenStudy (anonymous):

AHHHHHH and (0,1,0)!!

OpenStudy (amistre64):

the vectors we get from working the parametric form of the free variables become our Evectors

OpenStudy (anonymous):

Thanks amistre64!!!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

In a way, one has to solve \[\left( \begin{array}{ccc} 2 & 0 & -2 \\ 0 & 4 & 0 \\ -2 & 0 & 5 \\ \end{array} \right) \left( \begin{array}{c} x \\ y \\ z \\ \end{array} \right)=4 \left( \begin{array}{c} x \\ y \\ z \\ \end{array} \right) \]

OpenStudy (anonymous):

yep, got confused when i saw two zeros with my working out!

OpenStudy (anonymous):

You get x=0, z=0 and y can be anything different from 0

OpenStudy (anonymous):

what if x and z were not zeros?

OpenStudy (anonymous):

if x and y were not zero, 4 will not be an eigenvalue of the vector (x,y,z)

OpenStudy (anonymous):

assuming it was, and x had a value of 1 and z had a value of 2, y= 1, how would the matrix P and D be id P^-1 * M*P =D, where D is a diagonal matrix?

OpenStudy (anonymous):

The three eigenvectors are {{-1, 0, 2}, {0, 1, 0}, {2, 0, 1}}

OpenStudy (anonymous):

\[ P=\left( \begin{array}{ccc} -1 & 0 & 2 \\ 0 & 1 & 0 \\ 2 & 0 & 1 \\ \end{array} \right)\\ P^{-1} M P =\left( \begin{array}{ccc} 6 & 0 & 0 \\ 0 & 4 & 0 \\ 0 & 0 & 1 \\ \end{array} \right) \]

OpenStudy (anonymous):

yes but wouldnt that mean there would be four eigenvectors for the eigenvalues? what im trying to say is if you have a matrix M that has three eigenvalues but one of the eigenvalue has two eigen vectors? lets say one of the eigenvalue was 4 and we get M-(Lambda*I) = = 4 2 4 2 1 2 4 2 4 what would the eigen vector be?

OpenStudy (anonymous):

The maximum you can have is 3 independent eigenvectors.

OpenStudy (anonymous):

exactly!! so how would P look like? would there 5 columns? and would this mean that D would no more be a diagonal matrix?

OpenStudy (anonymous):

If Z is an eigenvector for 4 then 17 Z is an eigenvector for 4 then

OpenStudy (anonymous):

M (17 Z) = 17 M Z = 17(4 Z) = 4 ( 17 Z)

OpenStudy (anonymous):

P would look like \[ P=\left( \begin{array}{ccc} -1 & 0 & 2 \\ 0 & 17 & 0 \\ 2 & 0 & 1 \\ \end{array} \right) \] and it will still do the job. \[P^{-1} M P=\left( \begin{array}{ccc} 6 & 0 & 0 \\ 0 & 4 & 0 \\ 0 & 0 & 1 \\ \end{array} \right) \]

OpenStudy (anonymous):

Did you see what is going on?

OpenStudy (anonymous):

P Should be a 3x3 matrix whose rows are the independents eigenvectors.

OpenStudy (anonymous):

Usually we choose small numbers like 1 to make the computation easier.

OpenStudy (anonymous):

but what if an eigen value gives two eigenvectors?

OpenStudy (anonymous):

eigen space i think it was called

OpenStudy (amistre64):

as long as the results and up in linearly INdependant vectors, your good

OpenStudy (anonymous):

If they are independent we take both. In this case, the eigenvalue is a double roots.

OpenStudy (amistre64):

not all square matrixes are diagonalizable

OpenStudy (anonymous):

ahhhh okay!!

OpenStudy (anonymous):

thanks to all, well appreciated !

OpenStudy (anonymous):

yw